Please help me. I have tried these 3 on my own, but I am still not understanding

1) Two countries produce two goods and they each have 4 million hours of labor. Consider the table that shows the quantity produced per labor hour:

shirts corn
Alpha 4 12
Beta 2 10

If each country allocates 2 million hours of labor to each good then in total the two countries can together produce __ million shirts and ___ million brushels of corn.

2) Suppose that in a week an American worker can produce 60 shirts or 3 computers and a Chinese worker can produce 60 shirts or 1 computers. The country that specializes in computers will trade 1 computer in exchange for at least __ shirts. The country that buys the computer will pay not more than __ shirts. Enter whole numbers.

Consider the production possibilities frontier for a country that produces cars and wheat: 20 18 16 14 O 12 10 6 0 50100 50 200 250 300 350 400 Wheat (millions of bushels) In the absence of trade, if this country chooses to produce 6 million cars then they will also be able to produce wheat. Enter a whole number. Suppose the country specializes in wheat production and trades 100 million bushels of wheat in exchange for 8 million cars. They will be able to consume million bushels of million bushels of wheat. Enter a whole number.
Answer #1

Answer

In 1 hour alpha can produce 4 shirts or 12 bushels of corn. Hence if alpha uses 2 million hours each on production of corn and shirts then it can produce 4*2 million = 8 million shirts and 12*2 million = 24 million bushels of corn

In 1 hour Beta can produce 2 shirts or 10 bushels of corn. Hence if Beta uses 2 million hours each on production of corn and shirts then it can produce 2*2 million = 4 million shirts and 10*2 million = 20 million bushels of corn

Hence, Together they will produce 12 million shirts and 44 million bushels of corn

2) Suppose that in a week an American worker can produce 60 shirts or 3 computers. Hence Opportunity cost of producing 1 computer is 60/3 = 20 shirts. Chinese worker can produce 60 shirts or 1 computers. Hence Opportunity cost of producing 1 computer is 60 shirts. As, Hence Opportunity cost of producing 1 computer in china > Opportunity cost 1 computer in America.

Hence, America will specialize in Computer. Hence US will will trade 1 computer in exchange for at least 20 shirts. The country that buys the computer i.e. China will pay not more than 60 shirts.

3)

As we can see PPF is a straight line and using straight line equation we get the following equation of PPF

C = -(14/350)(W - 350)

where C and W are in millions

If C = 6 then W - 350 = -350*6/14 = -150

=> W = 200

Hence They will able to produce 200 million bushels of wheat

Suppose Country specialize in wheat then According to above PPF It will produce 350 million bushels of wheat and it has traded 100 million for 8 million Cars, Hence amount of wheat left = 350 million - 100 million = 250 million.

Hence It will able to consume 250 million bushels of Wheat